Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability: If 4 fair dice are thrown simultaneously

Expert replies
by ChessWriter » Sun Mar 11, 2012 11:37 am
If 4 fair dice are thrown simultaneously, what is the probability of getting atleast one pair?

A. 1/6
B. 5/18
C. 1/2
D. 2/3
E. 13/18

This is a question from the Veritas Prep Combinatorics and Probability book. I got this one wrong when I solved it. But, The book contains only the answer but not the detailed explanation for it.

I will post the Official Answer after some time
Join the discussion
Source: — Problem Solving |

by ChessWriter » Sun Mar 11, 2012 11:49 am
Immediately after I hit the "post" button after typing the above question, the method to solve this came to me. So, I am sharing my method and the official answer below.

One way to solve it is by considering the opposite - That on throwing the dice 4 times, you never get a pair.

In this case, the probability of what happens on the first throw does not matter. But the probability of getting a different number on the second, third and fourth throw are 5/6, 4/6 and 3/6 respectively.

So on multiplying we get 6/6*5/6*4/6*3/6 = 5/18

On substracting it from 1, we get 1-(5/18) = [spoiler]13/18[/spoiler]

Therefore [spoiler]E. 13/18[/spoiler] is the correct answer

But, I wonder, Is there any other way to solve this question?
Join the discussion

by pemdas » Sun Mar 11, 2012 2:46 pm
ChessWriter wrote:If 4 fair dice are thrown simultaneously, what is the probability of getting atleast one pair?

A. 1/6
B. 5/18
C. 1/2
D. 2/3
E. 13/18

This is a question from the Veritas Prep Combinatorics and Probability book. I got this one wrong when I solved it. But, The book contains only the answer but not the detailed explanation for it.

I will post the Official Answer after some time
total possible outcomes for dice 6*6*6*6 or 6^4
favorable outcomes or possible pairs =(4C2)*6=6^2=36

Thus, P(at least one pair)=(36+35+34+...1)/6^4, find mean of consecutive no-s 1...36, and calculate the sum -> P(at least one pair)= 666/6^4 = 6*111/6^4=111/216
Success doesn't come overnight!
Join the discussion